Please be thorough with the steps. The answers are in red. Consider the non -standard normal random variable Zua , where 5.5 and a 1.9 μ.σ ι P (ZJ.g<-5.69) P(Z<5.69) =0.4602 P (Zua> 3.
Answer #1

Answer)

As the data is normally distributed we can use standard normal z table to estimate the answers

Z = (x-mean)/s.d

Given mean = -5.5

S.d = 1.9

1)

P(x<-5.69)

Z = (-5.69 - (-5.5))/1.9 = -0.1

From z table, P(z<-0.1) = 0.4602

2)

P(x>-3.98)

Z = (-3.98 - (-5.5))/1.9 = 0.8

From z table, P(z>0.8) = 0.2119

3)

P(-5.69<x<-3.98) = p(x<-3.98) - p(x<-5.69)

P(x<-3.98)

Z = 0.8

From z table, p(z<0.8) = 0.7881

P(x<-5.69) = 0.4602

Required answer is 0.7881 - 0.4602 = 0.3280

4)

From z table, p(z<0.39) = 0.65

So, 0.39 = (a - (-5.5))/1.9

A = -4.7679

5)

Here we need to find

P(x-0.6 < u <x+0.6)

As s.d = 1

Z = 0.6

So answer would be

P(z<0.6) - p(z<-0.6) = 0.4515